# LIBERO-PRO: Towards Robust and Fair Evaluation of Vision-Language-Action Models Beyond Memorization

### v1

@@ Abstract
Figure 1: Model trajectory consistency under object and instruction modifications. Original task” denotes the standard evaluation, while “Replace,” “Change Pos.,” “Remove,” and “Messy instruction” introduce perturbations to the target object or instruction. Across all settings, the model yields nearly identical trajectories, suggesting a lack of genuine task understanding and environmental perception.
@@ 1 INTRODUCTION
Vision-Language-Action (VLA) models have recently emerged as a central paradigm for bridging perception, language understanding and action execution in embodied agentsSapkota et al. (2025). A crucial factor in advancing this line of research is the availability of standardized benchmarksJames et al. (2020); Liu et al. (2023); Mees et al. (2022); Li et al. (2024b); Nasiriany et al. (2024); Walke et al. (2023) that allow fair and reproducible comparison across different methods. Among them, LIBEROLiu et al. (2023) has rapidly become the most widely adopted evaluation suite for VLA, serving as the de facto standard for reporting performance. As shown in Table 1, models ranging from large-scale pre-trained to specialized systems are all benchmarked on LIBERO, making it the “common currency” of evaluation in this field. Consequently, the reliability of LIBERO as an evaluation protocol plays a decisive role in shaping research directions in VLA.
Figure 2: Success rates of OpenVLA, pi0, pi0.5, and univla under object position perturbations.
@@
However, the current LIBERO training–evaluation setup may suffer from serious limitations. In fact, the evaluation tasks are exactly the same as the training tasks, with the only difference being slight perturbations in the initial states of objects—perturbations that are almost invisible to the human eye. As a result, if a model overfits the training set, it can easily achieve high scores during evaluation, but its accuracy drops to nearly zero under minor changes in object initial positions or task instructions, as shown in Figure 2. Yet such results do not indicate that the model has learned to understand the environment or the task. On the contrary, we find that models rely on fixed mappings from the training tasks. As illustrated in Figure 1, when we introduced four types of modifications—(1) replacing the target object with an unrelated one; (2) relocating its initial position; (3) removing it entirely; and (4) corrupting the instruction with nonsensical tokens—their trajectories remained nearly identical. This outcome shows that, despite achieving over 90% on the standard LIBERO benchmark, the reported scores reflect rote recall rather than genuine task understanding. Obviously, the current evaluation protocol is inadequate and fails to effectively assess model architectures and algorithms.
@@
To address the above shortcomings, we systematically propose an extended evaluation suite, i.e., LIBERO-PRO, to enable a more comprehensive and reasonable assessment of VLA models. We design evaluation scenarios along four key dimensions: manipulated objects, initial states, task instructions, and environments. While ensuring that each task remains coherent and executable, we introduce carefully designed perturbations to the original LIBERO tasks to simulate real-world situations. Using this suite, we evaluate state-of-the-art models including OpenVLAKim et al. (2024) and pi0/0.5, and the results show that they exhibit near-complete failure under several of these settings: the models cannot cope with changes in object positions, nor can they correctly follow slightly rephrased task instructions—even when such instructions have appeared in other training contexts. These findings reveal that the commonly reported accuracies of over 90% are largely artifacts of task memorization, rather than evidence of genuine perception or task understanding. We therefore call on the community to adopt evaluation standards that incorporate such necessary perturbations, ensuring that reported progress reflects capabilities required for practical use. Our main contributions are as follows:
  • Through extensive experiments, we demonstrate that the current LIBERO evaluation protocol is fundamentally flawed: high scores primarily reflect rote memorization of training data rather than genuine task understanding or execution ability.
@@
  • We introduce LIBERO-PRO, a plug-and-play benchmark that incorporates carefully designed perturbations across four dimensions—manipulated objects, initial states, task instructions, and environments—and supports randomized combinations of these dimensions for more faithful evaluation.
  • We benchmark state-of-the-art models, including OpenVLA, pi0 and pi0.5, on LIBERO-PRO and show that they perform poorly, highlighting the inadequacy of the current LIBERO protocol and calling for LIBERO-PRO as a fairer and more reliable evaluation standard.
2 Related Work
Vision-Language-Action Models. Vision-Language-Action (VLA) models have rapidly emerged as a central paradigm for embodied AI, aiming to unify perception, natural language understanding, and action execution within a single framework. Recent years have witnessed an explosion of approaches, ranging from large-scale pre-trained models that leverage multimodal corpora to task-specific architectures tailored for robotic manipulation and navigationMa et al. (2025); Li et al. (2025b); Din et al. (2025). While these efforts have significantly advanced the field, they also underscore the urgent need for standardized evaluation benchmarksLiu et al. (2023); Mees et al. (2022); James et al. (2020). Without such benchmarks, it becomes difficult to conduct fair and reproducible comparisons across methods, hindering the cumulative progress of VLA research.
Benchmarks for VLA Evaluation. A number of benchmarks have been developed to enable systematic evaluation of VLA models, including RLBench James et al. (2020), CALVIN Mees et al. (2022), RoboCasa Nasiriany et al. (2024), and BridgeData Walke et al. (2023). These benchmarks provide valuable testbeds for task execution under vision and language guidance, yet their adoption has remained relatively limited within the broader community. In contrast, LIBERO Liu et al. (2023) has quickly emerged as the dominant benchmark and is now the most widely adopted protocol for evaluating VLA models. Its standardized task suite and unified reporting metrics have made it the de facto standard for performance comparison, to the extent that virtually all recent VLA studies report results on LIBERO. As such, LIBERO not only serves as the “common currency” of evaluation in this field, but also exerts a strong influence on research directions and claims of progress.
Table 1: Incomplete statistics of VLA models evaluated on LIBERO. Numbers shown to two decimal places. ‘–’ denotes not reported. Single: One checkpoint trained for all tasks. Multi: Separate checkpoint trained for each task.
Model Year Size Method Single/Multi Goal Object Spatial 10 Average
pi0.5 Intelligence et al. (2025) 2025 $\sim$7B SFT Single 0.98a 0.98 0.98a 0.92 0.97
CogACT Li et al. (2024a) 2024 $\sim$7B SFT - 0.90 0.98b 0.97b 0.89 0.98a
GeoVLA Sun et al. (2025) 2025 $\sim$7B SFT - 0.97b 0.99a 0.98a 0.97a 0.98a
OpenVLA Kim et al. (2024) 2024 $\sim$7B SFT Multi 0.98a 0.99a 0.98a 0.93b 0.97b
pi0 Black et al. (2024) 2024 $\sim$7B SFT Single 0.92 0.98b 0.97b 0.96b
UniVLA* Wang et al. (2025b) 2025 $\sim$8B SFT Single 0.94 0.99a 0.95 0.94b 0.96b
SwitchVLA Li et al. (2025c) 2025 SFT Single 0.93 0.93
TGRPO Chen et al. (2025) 2025 $\sim$7B RL Multi 0.91 0.91
GraspVLA Deng et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.82 0.89
UniVLA Bu et al. (2025) 2025 $\sim$7B SFT Multi 0.90 0.94 0.91 0.79 0.89
TriVLA Liu et al. (2025) 2025 $\sim$3B SFT - 0.90 0.94 0.91 0.73 0.87
MolmoAct Lee et al. (2025) 2025 $\sim$7B SFT Multi 0.88 0.95 0.87 0.77 0.87
CronusVLA Li et al. (2025a) 2025 $\sim$7B SFT Multi 0.91 0.95 0.90 0.69 0.86
SmolVLA Shukor et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.93 0.73 0.86
ThinkAct Huang et al. (2025) 2025 $\sim$7B SFT+RL - 0.87 0.91 0.88 0.71 0.84
World-VLA Cen et al. (2025) 2025 SFT Multi 0.83 0.96 0.88 0.60 0.82
Cot-VLA Zhao et al. (2025) 2025 $\sim$7B SFT - 0.88 0.92 0.88 0.69 0.81
VLA-RL Lu et al. (2025) 2025 $\sim$7B SFT+RL Multi 0.82 0.92 0.90 0.60 0.81
SpatialVLA Qu et al. (2025) 2025 $\sim$4B SFT Multi 0.79 0.90 0.88 0.56 0.78
Octo Team et al. (2024) 2024 $\sim$93M SFT Multi 0.85 0.86 0.79 0.51 0.75
VQ-VLA Wang et al. (2025c) 2025 $\sim$7B SFT Single 0.75 0.60 0.68
  • aBest (cell shaded light blue).   bSecond-best (cell shaded light green).
Limitations of Current Evaluation Practices. Several recent studies have noted that the high scores reported on LIBERO often fail to translate into reliable performance, even on tasks with only minor variations. This has raised serious concerns about the validity of its evaluation protocolZhou et al. (2025); Fang et al. (2025); Li (2025). In response, some researchers have explored stronger algorithms or proposed stress tests Gu et al. (2025); Guruprasad et al. (2024); Valle et al. (2025); Kube et al. (2025); Wang et al. (2025a)to encourage more robust VLA systems. However, the field still lacks a systematic and standardized framework for robust evaluation under LIBERO. The majority of works continue to rely on the original, flawed setup, reporting numbers that may not reflect genuine capability. To address this gap, we propose LIBERO-PRO, a plug-and-play extension of LIBERO designed to foster fair comparison across models and to advance research on building VLA systems that can withstand reasonable and practically necessary perturbations.
3 Problem Definition
3.1 Formal Task Definition
We define a VLA task to establish a rigorous basis for evaluating robotic systems.
$\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G),$ (1)
where $\ell$ is a natural language instruction specifying the task objective, $\mathcal{O}$ is a finite set of objects with attributes (e.g., shape, color, size, initial pose), $\mathcal{E}$ denotes the static environmental context (e.g., scene geometry, furniture layout), $p_{0}$ is a probability distribution over initial states $s_{0}$ (encoding init position and robot proprioception), and $G(s)$ is a binary goal predicate, returning $1$ if state $s$ satisfies $\ell$, and 0 otherwise.
The performance of a policy $\pi_{\theta}(a\mid s,\ell)$, parameterized by $\theta$, is measured by the expected success probability over a finite horizon $T$:
$\mathbb{P}_{s_{0}\sim p_{0},\pi_{\theta}}\left[G(s_{T})=1\right],$ (2)
where $s_{T}$ is the terminal state after $T$ steps. This formulation captures the integration of perception, language understanding, and action planning in VLA systems.
3.2 Current LIBERO Evaluation Protocol
The LIBERO benchmark, while valuable for VLA research, employs an evaluation protocol that may overestimate model competence due to its homogeneous train/test split. Each test task $\tau^{\prime}\in\mathcal{T}_{\mathrm{test}}$ mirrors a training task $\tau\in\mathcal{T}_{\mathrm{train}}$:
$\ell^{\prime}=\ell,\quad\mathcal{O}^{\prime}=\mathcal{O},\quad\mathcal{E}^{\prime}=\mathcal{E},\quad G^{\prime}=G,$ (3)
with the only difference being a minor perturbation in the initial state distribution:
$d_{\mathrm{TV}}(p_{0},p_{0}^{\prime})\leq\epsilon,\quad\epsilon\ll 1.$ (4)
Due to the small $\epsilon$, test tasks closely resemble training tasks, allowing policies to achieve high performance through memorization rather than robust generalization. This homogeneity may inflate performance metrics, masking limitations in compositional generalization and adaptability to diverse real-world scenarios, a critical concern for embodied AI.
Figure 3: LIBERO-PRO benchmark task overview. We extend the four original task categories in LIBERO by introducing object attribute perturbations, initial position perturbations, instruction perturbations, and environmental perturbations.
4 Methodology
4.1 Proposed Perturbation-Based Evaluation Framework
For a more reasonable evaluation of VLA models, we propose a framework that introduces controlled perturbations across four orthogonal dimensions, manipulated objects, initial states, task instructions, and environments. For a task $\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G)$, we generate perturbed tasks $\tau^{(k)}=\phi_{k}(\tau)$ using operators $\phi_{k}$, $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$, defined as follows:
Object Attribute Perturbation ($k=\mathrm{O}$): Modifies non-essential object attributes, such as color, texture, or size, while preserving semantic equivalence, e.g., changing a red cup to a blue cup. This tests the model’s robustness to superficial visual changes critical for real-world adaptability. The perturbed task is:
$\tau^{(\mathrm{O})}=\left(\ell,\mathcal{O}^{(\mathrm{O})},\mathcal{E},p_{0},G\right).$ (5)
Initial Position Perturbation ($k=\mathrm{I}$): Alters initial object positions to modify absolute and relative arrangements, ensuring physical plausibility, e.g., moving a cup relative to a plate. This probes the model’s spatial reasoning under varied layouts. The perturbed task is:
$\tau^{(\mathrm{I})}=\left(\ell,\mathcal{O},\mathcal{E},p_{0}^{(\mathrm{I})},G\right).$ (6)
Instruction Perturbation ($k=\mathrm{L}$): Perturbations are applied at both the semantic and task levels. Semantic perturbations adjust the linguistic expression while preserving the original task intent, such as paraphrasing “pick up the bottle” to “grab the bottle” or substituting with a jar. This evaluates the agent’s understanding of language and semantics. The perturbed task is:
$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O},\mathcal{E},p_{0},G\right),$ (7)
Task perturbations reasonably modify the task itself by replacing the target object or altering the action, e.g., changing “pick up the bottle” to “pick up the bowl.” We ensure that all actions and objects appear in the training set to support fair and meaningful generalization. The perturbed task is:
$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O}^{(\mathrm{L})},\mathcal{E},p_{0},G^{(\mathrm{L})}\right).$ (8)
Environment Perturbation ($k=\mathrm{E}$): Modifies the background’s appearance, e.g., changing a workspace, without affecting task feasibility, to assess robustness to environmental variations. The perturbed task is:
$\tau^{(\mathrm{E})}=\left(\ell,\mathcal{O},\mathcal{E}^{(\mathrm{E})},p,G\right).$ (9)
4.2 Perturbation Constraints
To ensure each perturbed task $\tau^{(k)}$ remains semantically valid, we impose two constraints. First, perturbations are confined within a dimension-specific neighborhood:
$x^{(k)}\in\mathcal{N}_{k}(x),\quad\mathrm{dist}_{k}(x,x^{(k)})\leq\delta_{k},$ (10)
where $x$ is the original component, $x^{(k)}$ is its perturbed version, and $\delta_{k}$ controls the magnitude of perturbation for dimension $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$. Second, we enforce non-trivial task-level separation:
$d_{\mathrm{TV}}(\tau,\tau^{(k)})>\epsilon,\quad\forall k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\},$ (11)
where $d_{\mathrm{TV}}$ is the total variation distance, ensuring that perturbations lead to substantive deviations. Multiple perturbed tasks are generated per category by applying $\phi_{k}$, with $\delta_{k}$ varied to characterize the sensitivity of policies under different perturbation magnitudes.
4.3 LIBERO-PRO Evaluation Benchmark
Building on these principles, we introduce LIBERO-PRO, a plug-and-play evaluation benchmark that enables systematic assessment under multiple dimensions of variation—object, position, instruction, task, and environment—as well as their arbitrary combinations, as shown in Figure 3. Specifically, for object variation, we create new assets for the four LIBERO tasks by modifying the appearance, size, and color of the original objects. For position variation, we define alternative spatial regions for all manipulable objects while strictly adhering to physical constraints and task definitions. For instruction variation, we provide three distinct paraphrased versions of each task description. For task variation, we redesign tasks in a principled manner to ensure feasibility, and supply new sets of relevant objects and target states for each modified task. For environment variation, we leverage the five environment categories built into the LIBERO benchmark and randomly substitute them across tasks. To avoid meaningless combinations, we restrict task variation from being combined with other types, while allowing all other dimensions of variation to be freely and randomly combined for more comprehensive evaluation.
5 Experiments
5.1 Settings
We conduct experiments on LIBERO-PRO, which applies controlled perturbations from four perspectives: object properties, initial positions, task instructions (covering both semantic and structural variations), and environments. We evaluate three representative models—OpenVLAKim et al. (2024), pi0Black et al. (2024), and pi0.5Intelligence et al. (2025). For fairness, OpenVLA is assessed using four official checkpoints trained on different LIBERO suites, whereas pi0 and pi0.5 are evaluated with a single official checkpoint. Consistent with the original LIBERO protocol, we set the number of evaluation episodes to 50 per task.
5.2 Main Results
We conducted a comprehensive evaluation of state-of-the-art vision-language-action (VLA) models Black et al. (2024); Kim et al. (2024) using the LIBERO-PRO evaluation suite, which introduces perturbations across four dimensions: objects, goals, spatial configurations, and task sets (see Tables 2, 3, 4, 5). Our findings are as follows: (1) Failure under simple perturbations. Despite achieving success rates above 90% on the standard LIBERO benchmark, models nearly collapse under changes to object positions or minor task modifications, even when constructed from training components. (2) Illusory robustness. Apparent stability against object and instruction perturbations is largely due to overfitting rather than genuine task understanding. (3) Model-specific sensitivity. Environmental perturbations affect models differently, depending on the degree of environmental divergence and reliance on memorized contexts. (4) Subtle capability differences. While most models fail under perturbations, Pi0.5 achieves a 0.38 success rate in the libero-goal task under position changes, compared to 0 for OpenVLA and Pi0, revealing differences masked by standard benchmark scores. Overall, these results show that current VLA models, despite excelling on the standard LIBERO, lack true robustness and generalization—highlighting the need for rigorous evaluation frameworks.
Table 2: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-goal benchmark. Accuracy on test (Ori: origina, Obj: object, Spa: spatial, Sem: semantic, Task: task-level, Env: environment). Action notation: $Open(x,y)$ = open target $y$ of container $x$; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Push(obj,loc)$ = push object objobj toward location locloc; $TurnOn(obj)$ = activate object objobj.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$Open(cabinet,drawer_{mid})$ 0.98 0.96 0.00 1.00 0.00 0.96 0.92 0.94 0.00 0.94 0.00 0.00 0.96 0.96 0.00 0.94 0.04 0.92
$Put(bowl,drawer_{top})$ 0.88 0.88 0.00 0.96 0.00 0.82 0.76 0.86 0.00 0.94 0.00 0.10 0.98 0.98 0.94 1.00 0.02 1.00
$Push(plate,stove_{front})$ 1.00 1.00 0.00 1.00 0.00 1.00 0.94 0.93 0.00 0.94 0.00 0.54 0.96 0.98 0.00 0.98 0.00 0.28
$Put(bowl,plate)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.99 0.00 0.96 0.00 0.96 0.90 0.90 0.00 0.92 0.02 0.04
$Put(bowl,stove)$ 0.94 1.00 0.00 1.00 0.00 1.00 0.98 0.97 0.00 0.98 0.00 0.98 0.98 0.96 0.00 0.98 0.04 0.78
$Put(bowl,cabinet_{top})$ 1.00 0.98 0.00 1.00 0.00 1.00 0.98 0.94 0.00 0.92 0.00 0.94 0.96 0.96 0.00 0.96 0.02 0.00
Put(cream_cheese,bowl)Put(cream\_cheese,bowl) 1.00 1.00 0.00 1.00 0.00 1.00 0.96 0.96 0.00 0.96 0.00 0.46 0.98 1.00 0.98 0.98 0.02 0.94
Put(wine_bottle,rack)Put(wine\_bottle,rack) 1.00 1.00 0.00 1.00 0.00 1.00 0.68 0.84 0.00 0.64 0.00 0.00 1.00 0.98 0.88 0.96 0.02 0.12
$Put(wine\_bottle,cabinet_{top})$ 0.96 0.83 0.00 0.88 0.00 0.88 0.94 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.98 0.98 0.02 0.48
$TurnOn(stove)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.00 0.96 0.00 0.00
Average 0.98 0.96 0.00 0.98 0.00 0.98 0.92 0.94 0.00 0.93 0.00 0.39 0.97 0.97 0.38 0.97 0.00 0.46
Table 3: Task-wise LIBER0 and LIBERO-PRO performance of three VLA models on the libero-spatial benchmark. Action notation: $Pick(src)$ = pick up object $bowl_{black}$ from location srcsrc and place $bowl_{black}$ onto/into target dstdst.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$Pick(between(plate,ramekin),plate)$ 1.00 0.95 0.00 0.88 0.00 0.02 1.00 0.97 0.00 0.94 0.00 1.00 1.00 0.96 0.02 0.98 0.00 0.66
Pick(table_center,plate)Pick(table\_center,plate) 1.00 0.98 0.00 1.00 0.00 1.00 1.00 1.00 0.00 0.98 0.00 0.98 1.00 1.00 0.00 0.98 0.02 0.36
$Pick(drawer_{top}(cabinet_{wood}),plate)$ 0.96 0.96 0.00 0.94 0.00 0.94 0.98 0.99 0.00 0.94 0.00 0.54 1.00 1.00 0.00 1.00 0.00 0.64
Pick(next_to(cookie_box),plate)Pick(next\_to(cookie\_box),plate) 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.94 0.00 1.00 0.00 1.00 1.00 0.98 0.00 1.00 0.02 0.86
Pick(next_to(plate),plate)Pick(next\_to(plate),plate) 1.00 0.95 0.00 0.94 0.00 0.92 0.96 1.00 0.00 0.94 0.00 0.84 0.90 0.92 0.00 0.90 0.00 0.22
Pick(next_to(ramekin),plate)Pick(next\_to(ramekin),plate) 1.00 0.99 0.00 1.00 0.00 0.98 1.00 0.80 0.00 1.00 0.00 0.48 1.00 0.96 0.12 1.00 0.02 0.00
Pick(on(cookie_box),plate)Pick(on(cookie\_box),plate) 1.00 0.99 0.00 0.98 0.00 1.00 0.98 0.92 0.00 0.98 0.00 0.96 1.00 1.00 0.00 0.98 0.00 1.00
$Pick(on(ramekin),plate)$ 0.92 0.95 0.00 0.96 0.00 1.00 0.90 1.00 0.00 0.96 0.00 0.10 1.00 1.00 0.98 1.00 0.02 0.00
$Pick(on(stove),plate)$ 1.00 0.99 0.00 1.00 0.00 1.00 0.96 0.99 0.00 0.96 0.00 0.04 0.96 0.96 0.02 0.94 0.00 0.82
$Pick(on(cabinet_{wood}),plate)$ 0.94 0.99 0.00 1.00 0.00 1.00 0.94 0.86 0.00 1.00 0.00 0.80 0.90 0.88 0.90 0.90 0.00 0.02
Average 0.98 0.97 0.00 0.97 0.00 0.89 0.97 0.95 0.00 0.97 0.00 0.60 0.98 0.97 0.20 0.97 0.01 0.46
Table 4: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-10 benchmark. Action notation: $TurnOn(obj)$ = activate object objobj; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Place(obj,loc)$ = place picked object objobj into/onto target locloc; $(obj).close$ = close container objobj.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$TurnOn(stove)\wedge Put(moka\_pot,stove)$ 1.00 0.95 0.00 1.00 0.00 0.00 0.74 0.67 0.00 0.74 0.00 0.00 0.92 0.90 0.18 0.90 0.00 0.94
$Put(bowl_{black},drawer_{bottom}(cabinet)).close$ 0.78 0.98 0.00 1.00 0.00 0.98 0.94 0.93 0.00 0.94 0.00 0.78 0.98 0.96 0.58 0.98 0.02 0.86
$Put(mug_{yellow\_white},microwave).close$ 1.00 0.98 0.00 0.94 0.00 0.96 0.88 0.74 0.00 0.88 0.00 0.00 0.98 0.98 0.00 1.00 0.00 0.46
$Put(\{moka\_pot_{1},moka\_pot_{2}\},stove)$ 0.92 0.64 0.00 0.92 0.00 0.78 0.22 0.28 0.00 0.22 0.00 0.00 0.98 0.98 0.00 0.98 0.02 0.26
$Put(\{alphabet\_soup,cream\_cheese\},basket)$ 0.96 0.98 0.00 1.00 0.00 1.00 1.00 0.96 0.00 1.00 0.00 0.16 1.00 0.98 0.00 0.98 0.00 0.52
$Put(\{alphabet\_soup,tomato\_sauce\},basket)$ 0.98 0.95 0.00 0.98 0.00 0.96 0.88 0.79 0.00 0.88 0.00 0.06 0.94 0.94 0.02 0.94 0.02 0.70
$Put(\{cream\_cheese,butter\},basket)$ 0.96 0.96 0.00 0.98 0.00 0.98 0.98 0.98 0.00 0.98 0.00 0.96 0.96 0.94 0.00 0.98 0.00 0.00
$Put((mug_{white},plate_{left})|(mug_{yellow\_white},plate_{right}))$ 0.84 0.85 0.00 0.92 0.00 0.98 0.86 0.72 0.00 0.86 0.00 0.00 1.00 0.98 0.00 1.00 0.02 0.22
$Put((mug_{white},plate)|(pudding_{choco},right\_of(plate)))$ 0.94 0.77 0.00 0.90 0.00 0.86 0.76 0.82 0.00 0.74 0.00 0.00 0.62 0.58 0.00 0.60 0.00 0.04
$Place(book,compartment_{back}(caddy))$ 1.00 1.00 0.00 0.98 0.00 1.00 0.94 0.98 0.00 0.94 0.00 0.76 0.94 0.94 0.00 0.94 0.00 0.56
Average 0.93 0.81 0.00 0.96 0.00 0.85 0.82 0.79 0.00 0.82 0.00 0.27 0.93 0.92 0.08 0.93 0.01 0.46
Table 5: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-object benchmark. Action notation: $Place(obj,loc)$ = pick up object objobj and place objobj into/onto target locloc.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
Place(alphabet_soup,basket)Place(alphabet\_soup,basket) 1.00 0.97 0.00 0.96 0.00 0.00 0.98 0.98 0.00 0.88 0.00 0.06 0.96 0.96 0.00 0.94 0.00 0.68
Place(bbq_sauce,basket)Place(bbq\_sauce,basket) 0.98 0.90 0.00 0.94 0.00 0.00 0.98 1.00 0.00 1.00 0.00 0.46 1.00 0.98 1.00 1.00 0.02 0.96
$Place(butter,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 1.00 0.98 0.00 1.00 0.00 0.00 0.96 0.98 0.54 0.98 0.00 0.00
Place(chocolate_pudding,basket)Place(chocolate\_pudding,basket) 0.98 1.00 0.00 0.98 0.00 0.00 0.98 0.96 0.00 1.00 0.00 0.18 1.00 0.96 0.00 0.94 0.02 0.82
Place(cream_cheese,basket)Place(cream\_cheese,basket) 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.10 1.00 0.00 0.76 0.98 1.00 0.00 0.94 0.00 0.88
$Place(ketchup,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 0.96 0.84 0.00 0.92 0.00 0.24 0.98 0.96 0.20 0.96 0.02 0.92
$Place(milk,basket)$ 1.00 0.98 0.00 0.96 0.00 0.00 1.00 0.84 0.00 1.00 0.00 0.16 1.00 1.00 0.00 0.98 0.00 0.80
Place(orange_juice,basket)Place(orange\_juice,basket) 1.00 1.00 0.00 1.00 0.00 0.00 0.92 1.00 0.00 1.00 0.00 0.62 0.98 1.00 0.00 0.96 0.02 0.28
Place(salad_dressing,basket)Place(salad\_dressing,basket) 0.94 1.00 0.00 1.00 0.00 0.00 0.98 0.98 0.10 1.00 0.00 0.00 1.00 0.98 0.00 0.98 0.00 1.00
Place(tomato_sauce,basket)Place(tomato\_sauce,basket) 0.96 1.00 0.00 1.00 0.00 0.00 0.98 0.82 0.00 0.92 0.00 0.48 0.98 0.98 0.00 0.96 0.00 0.92
Average 0.99 0.98 0.00 0.98 0.00 0.00 0.98 0.94 0.00 0.90 0.00 0.29 0.98 0.98 0.17 0.96 0.01 0.73
5.3 Does the Model Generalize to New Objects?

### v2

# Xueyang Zhou1Yangming Xu1Guiyao Tie1Yongchao Chen2Guowen Zhang1 Duanfeng Chu3 Pan Zhou1Lichao Sun4

@@ Abstract
Keywords Vision-Language-Action $\cdot$ Embodied AI $\cdot$ Benchmark $\cdot$ Robustness Evaluation
Figure 1: VLA models evaluated on LIBERO can be broadly grouped into five families according to their action decoding strategies: autoregressive, continuous, flow/diffusion, hybrid, and adapter-based methods. Although these different paradigms all achieve over 70% performance on LIBERO, their performance drops substantially on the more challenging LIBERO-PRO benchmark. This suggests that, despite strong results on LIBERO, current VLA models still suffer from limited robustness and generalization under more demanding evaluation settings.
@@ 1 INTRODUCTION
Vision-Language-Action (VLA) models have recently emerged as a fundamental paradigm for bridging perception, language understanding, and action execution in embodied agents [17]. A central catalyst for progress in this domain is the development of standardized benchmarks [7, 13, 15, 12, 16, 19], which enable reproducible, fair, and quantitative comparison across competing methods. Among these, LIBERO [13] has rapidly become the most widely adopted evaluation suite for VLAs, functioning as the de facto standard for performance reporting. As summarized in Figure 6, both large-scale pre-trained models and task-specific systems are benchmarked on LIBERO, rendering it the “common currency” for evaluation in this field. Consequently, the methodological rigor and reliability of LIBERO directly influence how research directions in VLA are defined and pursued.
@@
Despite its influence, the current LIBERO training–evaluation configuration exhibits fundamental limitations. Specifically, the evaluation tasks are identical to the training tasks, differing only by marginal perturbations in initial object states—variations so subtle as to be visually imperceptible. As a result, models that overfit to the training distribution can still achieve near-perfect evaluation scores, while their accuracy collapses under even minor changes in object position or task phrasing (Figure 6). Such inflated performance fails to reflect genuine task comprehension or environmental reasoning. Our analysis further reveals that VLA models often rely on memorized task mappings rather than learned representations. Specifically, when we applied four systematic modifications—(1) replacing the target object with an unrelated one, (2) relocating its initial position, (3) removing it entirely, and (4) corrupting the instruction with nonsensical tokens—the resulting trajectories remained nearly unchanged. These findings demonstrate that, despite achieving over 90% success on the standard LIBERO benchmark, current results primarily indicate rote recall rather than robust task understanding. Hence, the existing evaluation protocol is insufficient to rigorously assess the reasoning, compositionality, and generalization capacity of modern VLA architectures.
@@
To address these shortcomings, we introduce an extended evaluation suite, LIBERO-PRO, designed to provide a more comprehensive and reliable measure of VLA performance. LIBERO-PRO systematically constructs evaluation scenarios along four orthogonal dimensions—manipulated objects, initial states, task instructions, and environments—while ensuring that each perturbed task remains logically coherent and physically executable. Through carefully controlled perturbations, we simulate realistic variations that better approximate the diversity of real-world conditions. Using this suite, we evaluate state-of-the-art VLA models, including OpenVLA [8] and Pi0/Pi0.5, and observe near-complete performance collapse under several perturbation regimes: models fail to adapt to altered object positions and exhibit severe sensitivity to even minimal instruction paraphrasing, despite encountering semantically equivalent expressions during training. These findings reveal that the widely reported accuracies exceeding 90% largely reflect memorization rather than genuine perceptual grounding or task-level reasoning. We therefore advocate for evaluation protocols that incorporate structured perturbations, ensuring that reported progress aligns with the requirements of robust and generalizable embodied intelligence. Our key contributions are summarized as follows.
  • We empirically demonstrate that the current LIBERO evaluation protocol substantially overestimates model competence, as its high reported scores largely stem from memorization of training configurations rather than genuine understanding or transferable reasoning ability.
@@
  • We introduce LIBERO-PRO, a plug-and-play benchmark that integrates systematically designed perturbations across four core dimensions—manipulated objects, initial states, task instructions, and environments—supporting both single- and multi-dimensional randomized combinations for rigorous evaluation.
  • We benchmark leading VLA models, including OpenVLA, Pi0, and Pi0.5, on LIBERO-PRO and reveal significant performance degradation, underscoring the limitations of the original LIBERO protocol and motivating LIBERO-PRO as a fairer, more faithful, and reproducible evaluation standard.
2 Observing Overfitting Issues
Despite most VLA models achieving over 95% success on LIBERO, our qualitative experiments reveal several systematic failure patterns that collectively indicate severe overfitting to the training distribution. Specifically, we observe: (1) brittle visual perception, manifesting as object or location hallucinations; (2) insensitivity to instruction semantics, where actions are driven by trajectory recall rather than language grounding; and (3) a lack of compositionality, with models unable to integrate individually learned subtasks into coherent multi-step behaviors.
2.1 Visual perceptual hallucinations